# CHARME: A chain-based reinforcement learning approach for the minor embedding problem

### v1

@@ Abstract.
Quantum Annealing (QA) holds great potential for solving combinatorial optimization problems efficiently. However, the effectiveness of QA algorithms heavily relies on the embedding of problem instances, represented as logical graphs, into the quantum unit processing (QPU) whose topology is in form of a limited connectivity graph, known as the minor embedding Problem. Existing methods for the minor embedding problem suffer from scalability issues when confronted with larger problem sizes. In this paper, we propose a novel approach utilizing Reinforcement Learning (RL) techniques to address the minor embedding problem, named CHARME. CHARME  includes three key components: a Graph Neural Network (GNN) architecture for policy modeling, a state transition algorithm ensuring solution validity, and an order exploration strategy for effective training. Through comprehensive experiments on synthetic and real-world instances, we demonstrate that the efficiency of our proposed order exploration strategy as well as our proposed RL framework, CHARME. In details, CHARME  yields superior solutions compared to fast embedding methods such as Minorminer and ATOM. Moreover, our method surpasses the OCT-based approach, known for its slower runtime but high-quality solutions, in several cases. In addition, our proposed exploration enhances the efficiency of the training of the CHARME framework by providing better solutions compared to the greedy strategy.
copyright: acmlicensedjournalyear: XXXXdoi: XXXXXXX.XXXXXXXjournal: TQCjournalvolume: 37journalnumber: 4article: 111publicationmonth: 8ccs: Do Not Use This Code Generate the Correct Terms for Your Paperccs: Do Not Use This Code Generate the Correct Terms for Your Paperccs: Do Not Use This Code Generate the Correct Terms for Your Paperccs: Do Not Use This Code Generate the Correct Terms for Your Paper
@@ 1. Introduction
Quantum annealing (QA) is a quantum-based computational approach that leverage quantum phenomena such as entanglement and superposition to tackle complex optimization problems, in various domains such as machine learning (Pudenz and Lidar, 2013),(Caldeira et al., 2019),(Mott et al., 2017), bioinformatics (Mulligan and et al, 2019), (Fox and et al, 2022), (Ngo et al., 2023b), and networking (Guillaume and et al, 2022), (Ishizaki, 2019). QA tackles optimization problems through a three-step process. First, the problem is encoded using a logical graph which represents the quadratic unconstrained binary optimization (QUBO) formulation of a given problem. Next, the logical graph is embedded into a Quantum Processing Unit (QPU), which has its own graph representation known as hardware graph. The embedding aims to ensure that the logical graph can be obtained by contracting edges in the embedded subgraph of the hardware graph. This embedding process is commonly referred to as minor embedding. Finally, the quantum annealing process is repeatedly executed on the embedded QPU to find an optimal solution for the given problem.
@@
In this process, the minor embedding arises as a major bottleneck that prevents quantum annealing from scaling up. In particular, the topology of the hardware graph or its induced subgraphs may not perfectly align with that of the logical graph; therefore, minor embedding typically requires a significant number of additional qubits and their connections to represent the logical graph on the hardware graph. Consequently, this situation can lead to two primary issues that can degrade the efficiency of quantum annealing. First of all, the need for a significant number of additional qubits to embed the logical graph can strain the available resources of the hardware graph, leading to the scalability issues in term of limiting hardware resources. Furthermore, when the size of the logical graph increases, the running time required to identify a feasible embedding increases exponentially. As a result, quantum annealing with a prolonged embedding time can be inefficient in practical optimization problems, which require rapid decision-making.
@@
In the literature, there are two main approaches to address the above challenges in minor embedding, namely top-down and bottom-up. The top-down approach aims to find embeddings of complete graphs (Choi, 2011), (Klymko et al., 2012), (Boothby et al., 2016) in the hardware graph. While embedding a complete graph can work as a solution for any incomplete graph with the same or smaller size, the embedding process for incomplete graphs, especially sparse graphs, usually requires much fewer qubits compared to embedding a complete graph. As a result, the top-down approach may not be the most effective strategy for embedding sparse logical graphs. Although post-process techniques have been considered to mitigate this problem (Goodrich et al., 2018), (Serra et al., 2022), handling sparse graphs is a huge hindrance to the top-down approach.
@@
On the other hand, the bottom-up approach directly constructs solutions based on the topology of the logical and the hardware graphs. This approach involves computing minor embedding through either Integer Programming (IP) which finds an exact solution from predefined constraints (Bernal et al., 2020), or progressive heuristic methods which gradually map individual nodes of the logical graph to the hardware graph (Cai et al., 2014), (Pinilla and Wilton, 2019), (Ngo et al., 2023a). Compared to the top-down approaches, bottom-up methods are not constrained by the pre-defined embedding of complete graphs, allowing for more flexible embedding constructions. However, bottom-up approaches need to consider a set of complicated conditions associated to the minor embedding problem (which we will explain later). As a result, the computational cost required to obtain a feasible solution is substantial.
@@
In order to overcome the above shortcomings, we introduce a novel learning approach based on the Reinforcement Learning (RL). Leveraging its learning capabilities, RL can rapidly generate solutions for previously unseen problem instances using its well-trained policy, thus efficiently enhancing the runtime complexity. Furthermore, through interactions with the environment and receiving rewards or penalties based on the quality of solutions obtained from the environment, RL agents can autonomously refine their strategies over time. This enables them to explore state-of-the-art solutions, particularly beneficial for emerging and evolving challenges such as minor embedding. Thus, RL is commonly applied to solve optimization problems in quantum systems (Khairy et al., 2020; Chen et al., 2022; Wauters et al., 2020; Fan et al., 2022; Ostaszewski et al., 2021).
@@
Although RL has emerged as a powerful technique for handling optimization problems (Bengio et al., 2021), applying it to our minor embedding problem has encountered several challenges. The first challenge is the stringent feasibility requirement for solutions in the minor embedding problem. Specifically, feasible solutions must fulfill a set of conditions for the minor embedding problem including chain connection, global connection and one-to-many. These constraints lead to an overwhelming number of infeasible solutions, compared to the number of feasible ones within the search space. As a result, it is challenging for the RL framework to explore adequate number of feasible solutions during the training phase as well as the testing phase.
The second challenge lies in the policy design of the RL framework. In particular, the policy of the RL framework is a parameterized structure which receives the information derived from the current state as the input, and return the decision for the next state. In each step, the policy is queried to determine the next state. Thus, the efficiency of the RL framework heavily depends on the design of the policy which specifies the processing of state’s information. For example, a structure that extract uncorrelated relations from state’s information can hinder the convergence of the policy. In addition, the minor embedding problem considers two independent graphs which must be included in the state’s information. However, exploiting the relations between two graphs which are meaningful to the minor embedding problem is challenging. Therefore, there is a need for an efficient policy structure which can capture valuable relations in state’s information to facilitate the process of training policy’s parameters.
@@
To realize an RL-based approach for the minor embedding problem, we first introduce an initial design, named NaiveRL, which sequentially embeds each node in a given logical graph into an associated hardware graph. Based on this initial design, we further introduce CHARME, incorporating three key additional components, to tackle the aforementioned challenges. The first component is a GNN-based architecture, which integrates features from the logical graph, the hardware graph, and the embedding between the two graphs. This integration allows the architecture to return action probabilities and state values, serving as a model for the policy in the RL framework. The second component is a state transition algorithm which sequentially embed a node in the logical graph into a chain of nodes in the hardware graph in order to ensure the validity of the resulting solutions. Lastly, the third component is an order exploration strategy that navigates the RL agent towards areas in the search space with good solutions, facilitating the training process.
@@
Our proposed solutions are further evaluated extensively against three state-of-the art methods, including OCT-based, a top-down approach (Goodrich et al., 2018), and two bottom-up approaches Minorminer (Cai et al., 2014) and ATOM (Ngo et al., 2023a). The results illustrate that CHARME outperforms all state-of-the-art methods in term of qubit usage for very sparse logical graphs. In addition, the computational efficiency of CHARME  is comparable to that of the fastest method in terms of the running time.
@@
Organization. The rest of the paper is structured as follows. Section 2 introduces the definition of minor embedding and an overview of reinforcement learning. Our method and its theoretical analysis are described in Section 3. Section 4 presents our experimental results. Finally, section 5 concludes the paper.
@@ 2. Preliminaries
This section formally defines the minor embedding problem in the quantum annealing and presents the preliminaries needed for our proposed solutions. It includes a brief overview of ATOM for the minor embedding problem based on a concept of topology adaption (Ngo et al., 2023a), basic concepts of reinforcement learning, and Advantage Actor Critic (A2C) (Mnih et al., 2016), which will be partially adopted in our solution.
@@ 2.1. Minor Embedding Problem
Given an optimization problem with a set of binary variables $\mathbf{x}=\{x_{0},x_{1},...,x_{n}\}$ and quadratic coefficients $Q_{ij}$, the problem can be represented in a QUBO form as follows:
@@
The QUBO formulation is then encoded using a graph called logical graph. In the logical graph, each node corresponds to a binary variable $x_{i}$. For any two nodes $x_{i}$ and $x_{j}$, there exists an edge $(x_{i},x_{j})$ if the quadratic coefficient $Q_{ij}$ is nonzero.
The hardware graph is a representation of the topology of QPU with nodes corresponding to qubits and edges corresponding to qubits’ couplers. QA system of D-Wave consists of three QPU topologies, namely Chimera - the earliest topology, Pegasus - the latest topology, and Zephyr - next generation QPU topology. These topologies are all in a grid form: a grid of identical sets of nodes called unit cells. In this paper, we consider the Chimera topology. The reason is that embedding methods in Chimera topology may be translated to Pegasus topology without modification, because Chimera is a subgraph of Pegasus (Boothby et al., 2020). In other words, the method which we develop for Chimera model also works for Pegasus topology.
QUBO models a given optimization problem as a logical graph, and the topology of QPU used in QA as a hardware graph. In order to solve the final Hamiltonian which QA processes, we need to find an mapping from the logical graph to the hardware graph. Below we formally define the minor embedding problem:
Figure 1. A high-level overview of Minor Embedding problem in which QUBO formulation presented by a logical graph is embedded into a hardware graph.
@@
Given a logical graph $P$ and a hardware graph $H$, minor embedding problem seeks to find a mapping function $\phi:V_{P}\rightarrow\mathcal{P}(V_{H})$ satisfying three embedding constraints:
@@
    One-to-many: Two chains $\phi(u)$ and $\phi(v)$ in the hardware graph do not have any common nodes with $\forall u,v\in V_{P},u\neq v$.
We note that an embedding which satisfies the above three embedding constraints is referred as a feasible solution.
@@ 2.2. Heuristic approaches for the minor embedding problem
In this part, we provide an overview of heuristic solutions for the minor embedding problem, especially focused on ATOM (Ngo et al., 2023a).
Heuristic methods employ a step-by-step approach to construct solutions. In the work of Cai et al.(Cai et al., 2014), each step involves embedding a selected node from the logical graph into a suitable set of nodes in the hardware graph, while satisfying specific conditions. This iterative process continues until a feasible embedding is obtained. However, this method can encounter a problem when no appropriate set exists for the selected node. We call it isolated problem. To address this issue, the latest heuristic method, called ATOM(Ngo et al., 2023a), introduces an adaptive topology concept. Given a logical graph $P$ and a hardware graph $H$, this method consists of three main phases to find a feasible embedding from $P$ to $H$:
@@
    Initialization: A solution is initialized as $\phi\leftarrow\emptyset$, and an embedded set is initialized as $U_{P}\leftarrow\emptyset$. In addition, an permutation of nodes in the logical graph, denoted by $O=\{\bar{a}_{1},\dots,\bar{a}_{|V_{P}|}\}$, is precomputed heuristically. This permutation is consider as a node order to guide the embedding process, where each node in the logical graph is iteratively embedded into a chain of nodes in the hardware graph.
@@
    Path construction: In this phase, we iteratively embed nodes in $P$ into $H$ following the order of $O$ through $|V_{P}|$ steps. Specifically, in the step $t$, the node $\bar{a}_{t}$ is selected. Then, an algorithm, named NODE EMBEDDING, is designed to find a new embedding $\phi^{\prime}$ such that combination of $\phi^{\prime}$ and $\phi$ is a feasible embedding for $P[U_{P}\cup\{\bar{a}_{t}\}]$. If $\phi^{\prime}$ exists, $\phi$ is updated with $\phi^{\prime}$. This step is repeated for the subsequent $\bar{a}_{t}$ in the order $O$ until $O$ is empty. Otherwise, in cases where a feasible embedding $\phi^{\prime}$ cannot be obtained for node $\bar{a}_{t}$, it indicates the occurrence of an isolated problem. In such situations, phase 3 is triggered to handle this problem.
@@
ATOM method guarantees to return a feasible solution after at most $3|V_{P}|$ steps. In addition, NODE EMBEDDING and TOPOLOGY ADAPTING algorithms have low computational complexity. Thus, ATOM can achieve an efficient performance in practice. However, one limitation of this approach is that the heuristic strategy used to determine the embedding order $O$ may not be optimized, that can lead to a higher number of required qubits.

### v2

@@ Abstract.
Quantum annealing (QA) has great potential to solve combinatorial optimization problems efficiently. However, the effectiveness of QA algorithms is heavily based on the embedding of problem instances, represented as logical graphs, into the quantum processing unit (QPU) whose topology is in the form of a limited connectivity graph, known as the minor embedding problem. Because the minor embedding problem is an NP-hard problem (Goodrich et al., 2018), existing methods for the minor embedding problem suffer from scalability issues when faced with larger problem sizes. In this paper, we propose a novel approach utilizing Reinforcement Learning (RL) techniques to address the minor embedding problem, named CHARME. CHARME includes three key components: a Graph Neural Network (GNN) architecture for policy modeling, a state transition algorithm that ensures solution validity, and an order exploration strategy for effective training. Through comprehensive experiments on synthetic and real-world instances, we demonstrate the efficiency of our proposed order exploration strategy as well as our proposed RL framework, CHARME. In particular, CHARME yields superior solutions in terms of qubit usage compared to fast embedding methods such as Minorminer and ATOM. Moreover, our method surpasses the OCT-based approach, known for its slower runtime but high-quality solutions, in several cases. In addition, our proposed exploration enhances the efficiency of the training of the CHARME framework by providing better solutions compared to the greedy strategy.
copyright: acmlicensedjournalyear: XXXXdoi: XXXXXXX.XXXXXXXjournal: TQCjournalvolume: 37journalnumber: 4article: 111publicationmonth: 8ccs: Hardware Quantum computationccs: Computer systems organization Quantum computingccs: Computing methodologies Reinforcement learning
@@ 1. Introduction
Quantum annealing (QA) is a quantum-based computational approach that leverages quantum phenomena such as entanglement and superposition to tackle complex optimization problems, in various domains such as machine learning (ML) (Caldeira et al., 2019),(Mott et al., 2017), (Pudenz and Lidar, 2013), bioinformatics (Fox and et al, 2022), (Mulligan and et al, 2019), (Ngo et al., 2023b), and networking (Guillaume and et al, 2022), (Ishizaki, 2019). QA tackles optimization problems through a three-step process. First, the problem is encoded using a logical graph that represents the quadratic unconstrained binary optimization (QUBO) formulation of a given problem. Next, the logical graph is embedded into a Quantum Processing Unit (QPU), which has its own graph representation known as a hardware graph. The embedding aims to ensure that the logical graph can be obtained by contracting edges in the embedded subgraph of the hardware graph. This process is commonly referred to as minor embedding. Finally, the QA process is iteratively executed on the embedded QPU to find the optimal solution which results in the minimum value for the given QUBO.
@@
The minor embedding process is a major bottleneck that prevents QA from scaling. In particular, the topology of the hardware graph or its induced subgraphs may not perfectly align with that of the logical graph. Therefore, minor embedding typically requires a significant number of additional qubits and their connections to represent the logical graph on the hardware graph. Consequently, this situation can lead to two primary issues that can degrade the efficiency of QA. First, the need for a significant number of additional qubits to embed the logical graph can strain the available resources of the hardware graph, leading to scalability issues in terms of limiting hardware resources. Furthermore, when the size of the logical graph increases, the running time required to identify a feasible embedding increases exponentially. As a result, QA with a prolonged embedding time can be inefficient in practical optimization problems that require rapid decision-making.
@@
There are two main approaches to address the above challenges in minor embedding, namely top-down and bottom-up approaches. The top-down approach aims to find embeddings of complete graphs (Boothby et al., 2016), (Choi, 2011), (Klymko et al., 2012) in the hardware graph. Although embedding a complete graph can work as a solution for any incomplete graph with the same or smaller size, the embedding process for incomplete graphs, especially sparse graphs, usually requires much fewer qubits compared to embedding a complete graph. Specifically, experiments in (Goodrich et al., 2018) show that embedding techniques for sparse graphs can reduce the qubit usage by up to three times compared to the number of qubits required to embed complete graphs. As a result, the top-down approach may not be the most effective strategy for embedding sparse logical graphs. Although post-process techniques have been considered to mitigate this problem (Goodrich et al., 2018), (Serra et al., 2022), handling sparse graphs is a huge hindrance to the top-down approach.
@@
On the other hand, the bottom-up approach directly constructs solutions based on the topology of the logical and the hardware graphs. This approach involves computing a minor embedding through Integer Programming (IP) which finds an exact solution from predefined constraints (Bernal et al., 2020), or progressive heuristic methods which gradually map individual nodes of the logical graph to the hardware graph (Cai et al., 2014), (Ngo et al., 2023a), (Pinilla and Wilton, 2019). Compared to top-down approaches, bottom-up methods are not constrained by the predefined embedding of complete graphs, allowing for more flexible embedding constructions. However, bottom-up approaches need to consider a set of complicated conditions associated with the minor embedding problem. As a result, the computational cost required to obtain a feasible solution is substantial.
@@
To overcome the above shortcomings, we introduce a novel learning approach based on Reinforcement Learning (RL). Leveraging its learning capabilities, RL can rapidly generate solutions for previously unseen problem instances using its well-trained policy, thus efficiently enhancing the runtime complexity. Furthermore, through interactions with the environment and receiving rewards or penalties based on the quality of solutions obtained from the environment, RL agents can autonomously refine their strategies over time. This enables them to explore state-of-the-art solutions, which are particularly beneficial for emerging and evolving challenges such as minor embedding. Thus, RL is commonly applied to enhance the design of quantum systems (Khairy et al., 2020; Chen et al., 2022; Wauters et al., 2020; Fan et al., 2022; Ostaszewski et al., 2021; Saravanan and Saeed, 2024; LeCompte et al., 2023).
@@
Although RL cannot guarantee optimality, it has emerged as a powerful technique for handling optimization problems (Bengio et al., 2021). However, applying it to our minor embedding problem has exposed several challenges. The first challenge is the stringent feasibility requirement for solutions in the minor embedding problem. Specifically, feasible solutions must fulfill a set of conditions for the minor embedding problem, including chain connection, where each node in the logical graph $P$ is mapped to a connected subset of nodes (a ”chain”) in the hardware graph $H$; global connection, ensuring that if two nodes in $P$ are connected by an edge, at least one corresponding edge must exist between their mapped chains in $H$; and one-to-many, which requires that chains mapped from different nodes in $P$ do not overlap in $H$. All of these conditions are formally defined in Section 2. These constraints lead to an overwhelming number of infeasible solutions compared to the number of feasible solutions within the search space. As a result, it is challenging for the RL framework to explore a sufficient number of feasible solutions during the training phase as well as the testing phase.
The second challenge lies in the policy design of the RL framework. In particular, the policy of the RL framework is a mapping of observed input states to a choice of actions (Singh et al., 2000). As recent advances in deep learning in representing high-dimensional inputs, deep learning models are utilized to parameterize the policy function (Mnih et al., 2015). The deep learning model representing the policy (a.k.a. policy model) receives the information derived from the current state as input and returns the decision for the next action. Thus, the policy model is a key factor in the performance of the RL framework. In the context of applying RL to the minor embedding problem, the input state can include the topology information (of the given logical and hardware graphs) and the embedding information (which is a one-to-many mapping from the node set of the logical graph to the node set of the hardware graph). As the size of the two graphs increases, the number of state dimensions increases quadratically, potentially leading to an increase in the number of parameters required for the policy model to handle input states. On the other hand, reducing the state dimensions might degrade the policy model’s performance due to a lack of valuable features. Therefore, the policy model in the RL framework for the minor embedding problem must be capable of extracting and combining valuable features from both the topology and embedding information in order to scale well with the size of the logical and hardware graphs without compromising performance. However, designing such a policy model is challenging because the topology representations and the embedding information are not identical, complicating the data combination process.
@@
To realize an RL-based approach for the minor embedding problem, we first introduce an initial design, named NaiveRL, which sequentially embeds each node in a given logical graph into an associated hardware graph. Based on this initial design, we further introduce CHARME, incorporating three additional key components, to address the aforementioned challenges. The first component is a GCN-based architecture that integrates features of the logical graph, the hardware graph, and the embedding between the two graphs. This integration allows the architecture to return action probabilities and state values, serving as a model for the policy in the RL framework. The second component is a state transition algorithm that sequentially embeds a node in the logical graph into a chain of nodes in the hardware graph in order to ensure the validity of the resulting solutions. Lastly, the third component is an order exploration strategy that navigates the RL agent towards areas in the search space with good solutions, facilitating the training process.
@@
Our proposed solutions are extensively evaluated against three state-of-the-art methods, including an OCT-based technique, a top-down approach (Goodrich et al., 2018), and two bottom-up approaches Minorminer (Cai et al., 2014) and ATOM (Ngo et al., 2023a). The results illustrate that CHARME outperforms all state-of-the-art methods in terms of qubit usage for sparse logical graphs. For other graph types, while it does not outperform the OCT-based approach—which requires significant runtime for solution refinement—it remains the most efficient fast method, providing the best solutions within a short runtime. In addition, the computational efficiency of CHARME is comparable to that of the fastest method in terms of running time.
@@
Organization. The rest of the paper is structured as follows. Section 2 introduces the definition of minor embedding and an overview of RL. Our method and its theoretical analysis are described in Section 3. Section 4 presents our experimental results. Finally, Section 5 concludes the paper.
@@ 2. Preliminaries
This section formally defines the minor embedding problem in quantum annealing (QA) and presents the preliminaries needed for our proposed solutions. It includes a brief overview of ATOM for the minor embedding problem based on a concept of topology adaption (Ngo et al., 2023a), basic concepts of reinforcement learning (RL), and Advantage Actor Critic (A2C) (Mnih et al., 2016), which will be partially adopted in our solution.
@@ 2.1. Minor Embedding Problem
Given an optimization problem with a set of binary variables $\mathbf{x}=\{x_{1},x_{2},...,x_{n}\}$ and quadratic coefficients $Q_{ij}$, the problem can be represented in a quadratic unconstrained binary optimization (QUBO) form as follows:
@@
The QUBO formulation is then encoded using a logical graph. In the logical graph, each node corresponds to a binary variable $x_{i}$. For any two nodes $x_{i}$ and $x_{j}$, there exists an edge $(x_{i},x_{j})$ if the quadratic coefficient $Q_{ij}$ is non-zero.
The hardware graph is a representation of the Quantum Processing Unit (QPU) topology with nodes corresponding to qubits and edges corresponding to qubit couplers. The D-Wave QA system consists of three QPU topologies: Chimera (earliest), Pegasus (latest), and Zephyr (next generation). Each of these topologies is in a grid form: a grid of identical sets of nodes called unit cells. In this paper, we consider the Chimera topology because the embedding methods can be translated directly to the Pegasus topology without modification, since Chimera is a subgraph of Pegasus (Boothby et al., 2020). In other words, the method we develop for Chimera also works for Pegasus.
QUBO models a given optimization problem as a logical graph and the topology of a QPU used in QA as a hardware graph. In order to solve the final Hamiltonian that QA processes, we must find a mapping from the logical graph to the hardware graph. In the following, we formally define the minor embedding problem:
Figure 1. A high-level overview of the Minor Embedding problem where a QUBO formulation presented by a logical graph is embedded into a hardware graph.
@@
Given a logical graph $P$ and a hardware graph $H$, the minor embedding problem seeks to find a mapping function $\phi:V_{P}\rightarrow\mathcal{P}(V_{H})$ satisfying three embedding constraints:
@@
    One-to-Many: Two chains $\phi(u)$ and $\phi(v)$ in the hardware graph do not have any common nodes $\forall u,v\in V_{P},u\neq v$.
We note that an embedding which satisfies the above three embedding constraints is referred to as a feasible solution.
@@ 2.2. Heuristic approaches for the minor embedding problem
Heuristic methods employ a step-by-step approach to construct solutions. In the work of Cai et al. (Cai et al., 2014), each step involves embedding a selected node from the logical graph into a suitable set of nodes in the hardware graph, while satisfying specific conditions. This iterative process continues until a feasible embedding is obtained. However, this method can encounter a problem when there is no appropriate set for the selected node. We call this an isolated problem. To address this issue, the latest heuristic method, called ATOM(Ngo et al., 2023a), introduces the concept of adaptive topology. Given a logical graph $P$ and a hardware graph $H$, this method consists of three main phases to find a feasible embedding from $P$ to $H$:
@@
    Initialization: A solution is initialized as $\phi\leftarrow\emptyset$, and an embedded set is initialized as $U_{P}\leftarrow\emptyset$. In addition, a permutation of the nodes in the logical graph, denoted by $O=(\bar{a}_{1},\dots,\bar{a}_{|V_{P}|})$, is precomputed heuristically. This permutation is considered as a node order to guide the embedding process, where each node in the logical graph is iteratively embedded into a chain of nodes in the hardware graph.
@@
    Path Construction: In this phase, we iteratively embed nodes from $P$ into $H$ following the order of $O$ through $|V_{P}|$ steps. Specifically, in step $t$, the node $\bar{a}_{t}$ is selected. Then, an algorithm, named Node Embedding, is designed to find a new embedding $\phi^{\prime}$ such that the combination of $\phi^{\prime}$ and $\phi$ is a feasible embedding for $P[U_{P}\cup\{\bar{a}_{t}\}]$. If $\phi^{\prime}$ exists, $\phi$ is updated with $\phi^{\prime}$. This step is repeated for the subsequent $\bar{a}_{t}$ in the order $O$ until $O$ is empty. Otherwise, in the cases where a feasible embedding $\phi^{\prime}$ cannot be obtained for node $\bar{a}_{t}$, it indicates the occurrence of an isolated problem. In such situations, phase 3 is triggered to handle this problem.
@@
The ATOM method guarantees to return a feasible solution after at most $3|V_{P}|$ iterations of phase 2. In addition, denoting the average degree of the logical graph as $d$, the Node Embedding and Topology Adapting algorithms have polynomial complexity, with $O(|V_{H}|\times d)$ and $O(|V_{H}|)$, respectively (Ngo et al., 2023a). Thus, ATOM has a total complexity of $O(|V_{P}|\times|V_{H}|\times d)$, making it efficient for hardware graphs in practical QA machines, which have qubit limits (i.e., $V_{H}$) ranging from $5000$ to $7000$ (Vert et al., 2024). However, one limitation of this approach is that the heuristic strategy used to determine the embedding order $O$ may lead to a high number of required qubits. With the use of RL, CHARME, is able to learn and provide better orders. Additionally, we incorporate Node Embedding and Topology Adapting as two subroutines in our method.